A factory produces a closed rectangular parallelepiped vats having the capacity of 10 cubic meters. Find the dimensions that will make the cost of the lining a minimum?

1
Expert's answer
2022-01-20T16:47:25-0500

The volume of a rectangular parallelepiped is "xyz" which is 10 cubic meters


"xyz=10"

The surface area of the figure is "2xy+2xz+2yz" which is to be minimized.

Using Langrage method.


"F(x,y,z,\lambda)=(2xy+2xz+2yz)+\lambda(xyz-10)\\\\
\frac{\partial F}{\partial x}=2y+2z+\lambda yz\\\\
\frac{\partial F}{\partial y}=2x+2z+\lambda xz\\\\
\frac{\partial F}{\partial z}=2x+2y+\lambda xy\\
\frac{\partial F}{\partial \lambda}=xyz-10\\"


Equating the partial derivatives to zero.


"2y+2z+\lambda yz=0~~~~~~~~~~~~~~~~~~~(1)\\
2x+2z+\lambda xz=0~~~~~~~~~~~~~~~~~~~(2)\\
2x+2y+\lambda xy=0~~~~~~~~~~~~~~~~~~~(3)\\
xyz-10=0~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(4)"


Solving (1) and (2) together, we have that;


"x=y"


Also, solving (2) and (3) together, we have that;


"y=z"


Thus,


"x=y=z"


Substituting this into (4), we have that;


"x^3-10=0\\
x^3=10\\
x=\sqrt[3]{10}"

Hence the dimension to minimize the lining is


"\sqrt[3]{10}\times \sqrt[3]{10}\times \sqrt[3]{10}"